# The Unseen Frontier: Pushing the Limits of LLM Sparsity with Surrogate-Free ADMM

### v1

@@ 1 Introduction
Large language models (LLMs) have become indispensable tools across various fields, from creative industries to scientific research, but their immense size incurs a tremendous amount of memory, computation, and energy consumption, posing a significant challenge to their widespread deployment (Kaplan et al., 2020; Bommasani, 2021; Faiz et al., 2024). Neural network pruning can offer a viable solution to this problem by removing redundant parameters without compromising performance (LeCun et al., 1989; Han et al., 2015; Hoefler et al., 2021). Indeed, the research community has responded to this challenge with a surge of innovative methodologies, demonstrating that LLMs can be made more compact and efficient through effective pruning techniques (Frantar & Alistarh, 2023; Sun et al., 2024; Boža, 2024; Meng et al., 2024; Fang et al., 2024; Liu et al., 2025; Lee et al., 2025).
@@
[…]
@@ 2 Problem statement
The long-standing research of neural network pruning, aimed at enhancing the efficiency of large models (LeCun et al., 1989; Han et al., 2015), has recently made significant progress in its application to LLMs (Frantar & Alistarh, 2023; Sun et al., 2024; Boža, 2024; Liu et al., 2025). While effective, these methods decline sharply and fail to maintain performance beyond a moderate level of sparsity around 50-60%. For example, the recent study of Zhang et al. (2024) to evaluate these methods report that their performance begins to collapse after 70% sparsity. This deterioration is also evident in other recent works that, notwithstanding the relative advantage over existing methods, the majority still suffer from severely degraded performance in high-sparsity regimes, with perplexity often increased more than an order of magnitude (Boža, 2024; Meng et al., 2024). In fact, this stands in stark contrast to historical precedents, where extreme sparsity of say 90% or higher was commonly achieved (Frankle & Carbin, 2019; Lee et al., 2019). Consequently, researchers has begun to theorize the underlying causes, attributing the failure to compounding layer-wise errors and the explosion of reconstruction error (Shin et al., 2024; Huang et al., 2025).
@@
First of all, it is hard to solve (2) exactly without errors, in other words, the distance (3) cannot be zero realistically. This is due to the high cost of exactly solving sparse linear regression (Natarajan, 1995). In fact, this leads to layer-wise solvers relying on saliency-based heuristics to find approximate solutions (Frantar & Alistarh, 2023; Sun et al., 2024; Meng et al., 2024). Without zero layer-wise reconstruction errors, even small errors from each layer can compound into large overall errors, which has been observed to pose non-trivial harm to performance (Shin et al., 2024; Huang et al., 2025).
@@ 5 Experiments
We present a series of concrete experiments to validate Elsa in this section. Specifically, we show that Elsa (i) effectively prunes models to extreme high sparsity levels across a wide range of models and scales (Section˜5.1), (ii) scales efficiently to large models up to 27B (Section˜5.2), and (iii) adapts to other sparsity patterns such as N:M semi-structured sparsity or non-uniform sparsity found by evolutionary strategies (Section˜5.3). We also provide an ablation study on the choice of objective functions and generalized projection (Section˜5.4).
We compare Elsa to the following methods: Magnitude (Han et al., 2015), SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), ALPS (Meng et al., 2024), L-ADMM (Layer-wise ADMM) (Boža, 2024), SAFE (Lee et al., 2025), and SparseLLM (Bai et al., 2024). These methods are applied to four different architectures including OPT (Zhang et al., 2022), Gemma-2 (Team et al., 2024), and LLaMA-2/3 (Touvron et al., 2023; Grattafiori et al., 2024) across a wide range of scales from 125M to 27B. We report perplexity and zero-shot prediction accuracy of pruned models. All experiment settings can be found in Appendix˜B. The source code to reproduce results will be made available at https://github.com/log-postech/elsa.
@@ 5.1 Main results
This extends to downstream task performance, as shown in Figure˜4. Each radar plot reports per-task accuracy at high sparsity (70–90%), with the enclosed area reflecting the average accuracy across tasks. At 70% sparsity, Elsa is competitive with leading methods, but a clear gap emerges as sparsity increases. From 70% to 80% sparsity, other methods lose 10–20%p accuracy on tasks such as Winogrande and ARC-E, while Elsa degrades by less than half as much. At 90%, most methods collapse, whereas Elsa retains the highest accuracy on 6 out of 7 tasks, with an average margin of 6.06%p. This demonstrates that Elsa maintains generalization far better than existing methods at high sparsity.
We believe that these results collectively establish the effectiveness of Elsa for extreme sparsity.
Figure 4: Zero-shot accuracy of pruned LLaMA-2-7B models. Elsa outperforms other methods for most tasks, with the performance gap widening as sparsity increases, highlighting its strong generalization capability. Full numerical results are provided in Table˜7 of Appendix C.
@@ 5.2 Scaling to large-r models
Figure 5: Perplexity of Gemma-2-27B. Elsa-$_{\text{-L}}$ achieves the lowest perplexity, confirming its strength.
To further validate the scalability of our principle, we apply Elsa-$_{\text{-L}}$ to 27B-scale (Gemma-2-27B). Specifically, we employ the low-precision optimizer adam8bit (Dettmers et al., 2022) for $x$-update Equation˜7, and store the auxiliary variables $(u,z)$ in (bf16 ,fp8) precision (Kalamkar et al., 2019; Micikevicius et al., 2022). To save the states in low-precision, we apply quantization function $\mathcal{Q}$ at the tensor level with dynamically updated scales $s^{t}$ after each update. This design reduces the memory footprint by 66% compared to Elsa, enabling pruning at 27B scale under limited resources. We conduct this experiment at 90% sparsity level to test whether Elsa-$_{\text{-L}}$ can maintain performance under extreme compression. As shown in Figure˜5, Elsa-$_{\text{-L}}$ achieves the lowest perplexity among all compared methods (with some omitted due to infeasible memory requirements), outperforming the strongest competing method by a factor of $4\times$. These results reinforce our main finding that Elsa preserves model quality even at extreme sparsity and scale. Additional implementation details can be found in Section˜B.3.
5.3 Other sparsity patterns
Perplexity ($\downarrow$) Tasks ($\uparrow$)
0% Dense 5.47 7.26 43.35 76.26 77.68 57.14 31.40 62.82 69.06 59.67
Wanda 12.13 15.63 30.46 61.83 68.26 41.28 24.20 53.07 62.51 48.80
SparseGPT 10.87 13.61 30.97 64.06 67.61 43.47 24.20 56.32 66.38 50.43
L-ADMM 10.19 12.51 32.85 66.04 68.81 45.05 25.40 56.32 66.38 51.55
ALPS 9.945 12.09 34.47 68.86 73.79 49.40 27.60 55.60 67.25 53.85
SAFE 9.914 12.53 30.46 63.43 66.42 44.66 21.60 53.07 61.80 48.78
SparseLLM 11.29 13.95 30.55 61.91 71.10 43.62 24.40 57.40 65.82 50.69
Elsa 10.15 12.34 31.49 61.24 66.36 47.87 23.60 52.71 63.85 49.59
Wanda 8.603 11.33 34.47 67.05 72.87 46.98 26.80 54.15 66.93 52.75
SparseGPT 8.508 10.81 34.81 68.56 71.77 48.26 27.80 56.68 68.11 53.71
L-ADMM 8.12 10.37 35.58 68.18 72.48 49.45 28.80 58.12 67.17 54.25
ALPS 8.103 10.29 33.28 65.19 68.75 45.96 26.20 55.96 65.98 51.62
SAFE 8.043 10.47 31.57 66.84 68.04 48.55 23.40 53.07 65.04 50.93
SparseLLM 8.679 11.04 34.90 68.35 75.14 48.28 26.20 56.68 66.46 53.71
Elsa 9.20 11.47 32.25 64.69 69.42 49.90 27.40 53.07 63.22 51.42
We first evaluate Elsa for its adaptivity to N:M semi-structured sparsity, a setting designed for some current hardwares to accelerate computations (Sun et al., 2024; Fang et al., 2024). The results of both perplexity and zero-shot prediction accuracy are reported in Table˜1. Elsa is roughly on par with existing methods, and yet, it is noteworthy that these 2:4 and 4:8 sparsity patterns only ensure 50% sparsity. More importantly, these results indicate that Elsa can easily adapt to arbitrary constraints of moderate sparsity levels without much trouble.
Method Wiki()($\downarrow$) C4()($\downarrow$)
SparseGPT 85.84 98.35
OWL 48.07 52.32
EvoPress 28.76 33.72
Elsa (EvoPress) 26.11 29.33
Elsa 24.97 29.09
We also compare Elsa with non-uniform sparsity allocation based pruning methods. Specifically, we compare to OWL (Yin et al., 2024a) that allocates sparsity based on outlier distributions and to EvoPress (Sieberling et al., 2024) that uses an evolutionary search strategy to determine the non-uniform sparsity levels over different layers. We further set up a method that overrides Elsa with the mask found by the evolutionary strategy of EvoPress. Note that the sparsity level is set to be 70%; it is simply because these methods only works or reports up to this level. The results are presented in Table˜2. One can see that Elsa substantially outperforms OWL and shows an improvement over EvoPress as well: to elaborate, for instance, it achieves the C4 perplexity of $29.09$, compared to $33.72$ for EvoPress and $52.32$ for OWL. Notably, adopting the non-uniform mask found by EvoPress within Elsa yields some gains over the EvoPress itself, but it still falls short of the uniform allocation in Elsa, demonstrating the strength of our surrogate-free global formulation.
5.4 Ablations
Figure 6: Effect of NTP on data efficiency and perplexity.
Table 3: Effectiveness of geometric projection (✓).
Sparsity
70% 29.44 28.24
80% 40.06 37.50
90% 65.41 48.69
We also evaluate the effectiveness of the objective-aware projection on high-sparsity regimes. Specifically, we measure the perplexity of LLaMA-3.2-3B model pruned for 70-90% sparsity levels by turning on and off of the projection and report the results in Table˜3. The benefit of objective-aware projection grows with sparsity: perplexity gap increases from 1.20 at 70% sparsity to 2.56 at 80%, and widens further at 90%. This demonstrates that incorporating objective-aware importance into the projection step can be beneficial particularly in high sparsity regimes.
@@ 6 Discussion
In this work, we confront the problem of moderate sparsity in LLMs through a critical inspection into the current practice, revealing that the prevailing reliance on the sequential layer-wise reconstruction surrogate may have been constraining the path toward more extreme sparsities. This led us to develop Elsa and Elsa-$_{\text{-L}}$, enabling us to push the sparsity from 50-70% up to 80-90% while maintaining strong language modeling performance. Grounding on optimization principles ensures that our principle effectively solves the true LLM objective as is, while also facilitating the development of advanced techniques that are both theoretically sound and effective for sparsifying LLMs, which we believe were instrumental in attaining strong practical results.
@@
There are many promising directions to pursue for future work: (i) alternative efficiency strategies through advanced memory-efficient and derivative-free optimizers, (ii) system-level advancements in memory offloading, and (iii) extensions to advanced architecture such as Mixture-of-Experts (Mu & Lin, 2025) and multi-modal large language models (Yin et al., 2024b). To conclude, our work validates that the frontier of LLM sparsity can still be expanded by offering a concrete strategy supported by strong empirical evidence. We hope it sets the stage for future breakthroughs and innovations in new directions that have thus far received relatively limited attention.
@@ Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)), the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) ( RS-2023-00210466, RS-2023-00265444, RS-2025-02264052).
@@
Batch size 8
Training steps 4096
@@
Sparsity OPT-125M OPT-1.3B Gemma-2-2B LLaMA-3.2-3B LLaMA-2-7B LLaMA-2-13B
$\eta$ 50% 1e-5 1e-1 2e-1 1e-1 1e-1
60% 5e-5 1e-2 1e-2 2e-1
70% 1e-4 5e-3 1e-2 5e-3 2e-2
80% 5e-3 2e-2 5e-2
90% 2e-4 1e-3 5e-4 1e-3 1e-3 2e-3
$\lambda$ 50% 1e-2
60% 5e-3
70% 2e-3 5e-5 5e-5
@@
90% 1e-3 1e-4 5e-5 1e-4 1e-4 5e-5
$\lambda$ schedule constant cosine constant cosine

### v2

@@ 1 Introduction
Large language models (LLMs) have become indispensable tools across various fields, from creative industries to scientific research, but their immense size incurs a tremendous amount of memory, computation, and energy consumption, posing a significant challenge to their widespread deployment (Kaplan et al., 2020; Bommasani, 2021; Faiz et al., 2024). Neural network pruning can offer a viable solution to this problem by removing redundant parameters without compromising performance (LeCun et al., 1989; Han et al., 2015; Hoefler et al., 2021). Indeed, the research community has responded to this challenge with a surge of innovative methodologies, demonstrating that LLMs can be made more compact and efficient through effective pruning techniques (Frantar and Alistarh, 2023; Sun et al., 2024; Boža, 2024; Meng et al., 2024; Fang et al., 2024; Liu et al., 2025; Lee et al., 2025).
@@
We show that our approach can be applied to a wide range of LLM models and scales from 125M to 13B number of parameters. Across this range, Elsa remains stable in the highly sparse regime, whereas competing methods frequently collapse with order-of-magnitude perplexity blow-ups. Importantly, these gains translate into practical benefits: we obtain up to $\times 7.80$ memory reduction and $\times 3.98$ inference speedup without degrading model usability. We provide a flexible implementation as well, which incorporates memory-efficient designs including quantized optimizer states and enables pruning even for 27B-parameter models with 55% lower memory footprint, demonstrating extended potential at scale. Based on classic optimization theory, we also provide a convergence guarantee for our solver to ensure theoretical soundness alongside empirical findings.
@@ 2 Problem statement
The long-standing research of neural network pruning, aimed at enhancing the efficiency of large models (LeCun et al., 1989; Han et al., 2015), has recently made significant progress in its application to LLMs (Frantar and Alistarh, 2023; Sun et al., 2024; Boža, 2024; Liu et al., 2025). While effective, these methods decline sharply and fail to maintain performance beyond a moderate level of sparsity around 50-60%. For example, the recent study of Zhang et al. (2024) to evaluate these methods report that their performance begins to collapse after 70% sparsity. This deterioration is also evident in other recent works that, notwithstanding the relative advantage over existing methods, the majority still suffer from severely degraded performance in high-sparsity regimes, with perplexity often increased more than an order of magnitude (Boža, 2024; Meng et al., 2024). In fact, this stands in stark contrast to historical precedents, where extreme sparsity of say 90% or higher was commonly achieved (Frankle and Carbin, 2019; Lee et al., 2019). Consequently, researchers has begun to theorize the underlying causes, attributing the failure to compounding layer-wise errors and the explosion of reconstruction error (Shin et al., 2024; Huang et al., 2025).
@@
First of all, it is hard to solve (2) exactly without errors, in other words, the distance (3) cannot be zero realistically. This is due to the high cost of exactly solving sparse linear regression (Natarajan, 1995). In fact, this leads to layer-wise solvers relying on saliency-based heuristics to find approximate solutions (Frantar and Alistarh, 2023; Sun et al., 2024; Meng et al., 2024). Without zero layer-wise reconstruction errors, even small errors from each layer can compound into large overall errors, which has been observed to pose non-trivial harm to performance (Shin et al., 2024; Huang et al., 2025).
@@ 5 Experiments
We present a series of concrete experiments to validate Elsa in this section. Specifically, we show that Elsa (i) effectively prunes models to extreme high sparsity levels across a wide range of models and scales (Section˜5.1), (ii) can further push the limits to extreme sparsity (up to 99%) (Section˜5.2), (iii) can accelerate inference while reducing memory requirements (Section˜5.3), and (iv) scales efficiently to large models up to 27B (Section˜5.4). Lastly, we analyze the cost of Elsa compared to existing methods at (Section˜5.5). We also provide extension to other sparsity patterns such as N:M semi-structured sparsity or non-uniform sparsity, and an ablation study on the choice of objective functions and generalized projection on Appendix˜C.
We compare Elsa to the following methods: Magnitude (Han et al., 2015), SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024), ALPS (Meng et al., 2024), L-ADMM (Layer-wise ADMM) (Boža, 2024), SAFE (Lee et al., 2025), and SparseLLM (Bai et al., 2024). These methods are applied to four different architectures including OPT (Zhang et al., 2022), Gemma-2 (Team et al., 2024), and LLaMA-2/3 (Touvron et al., 2023; Grattafiori et al., 2024) across a wide range of scales from 125M to 27B. We report perplexity and zero-shot prediction accuracy of pruned models. All experiment settings can be found in Appendix˜B. The source code to reproduce results will be made available at https://github.com/log-postech/elsa.
@@ 5.1 Main results
This extends to downstream task performance, as shown in Figure˜4. Each radar plot reports per-task accuracy at high sparsity (70–90%), with the enclosed area reflecting the average accuracy across tasks. At 70% sparsity, Elsa is competitive with leading methods, but a clear gap emerges as sparsity increases. From 70% to 80% sparsity, other methods lose 10–20%p accuracy on tasks such as Winogrande and ARC-E, while Elsa degrades by less than half as much. At 90%, most methods collapse, whereas Elsa retains the highest accuracy on 6 out of 7 tasks, with an average margin of 6.06%p. This demonstrates that Elsa maintains generalization far better than existing methods at high sparsity. We believe that these results collectively establish the effectiveness of Elsa for high sparsity.
Figure 4: Zero-shot accuracy of pruned LLaMA-2-7B models. Elsa outperforms other methods for most tasks, with the performance gap widening as sparsity increases, highlighting its strong generalization capability. Full numerical results are provided in Table˜11 of Appendix C.
Table 1: Memory savings and inference accelerations of Elsa on LLaMA-2-7B.
Dense 50% 70% 90% 95%
Latency (s) 1.84 1.37 ($\times$1.34) 0.95 ($\times$1.94) 0.72 ($\times$2.50) 0.46 ($\times$4.00)
Tokens/s 54.47 72.81 ($\times$1.33) 104.89 ($\times$1.93) 139.47 ($\times$2.56) 216.98 ($\times$3.98)
Memory (MB) 13596 8870 ($\times$1.53) 5603 ($\times$2.42) 2918 ($\times$4.60) 1743 ($\times$7.80)
Table 2: Perplexity on Wiki/C4 at extreme sparsity.
Sparsity Method Wiki C4
0.90 Wanda + LoRA 92.66 65.56
Wanda + Full 42.40 34.87
Elsa 26.97 23.14
0.95 Wanda + LoRA 371.0 143.0
Wanda + Full 84.30 53.62
Elsa 38.91 28.39
0.99 Wanda + LoRA 588.3 247.5
Wanda + Full 146.37 71.64
Elsa 55.94 40.10
5.2 Towards extreme sparsity
To assess whether Elsa remains effective under extreme sparsity, we further evaluate LLaMA 2-7B at 95% and 99% sparsity. We additionally consider retraining after pruning with Wanda (Sun et al., 2024) for the baseline, and the experimental details can be found at Section˜B.3.
As shown in Table 2, Elsa consistently achieves substantially lower perplexity on both WikiText and C4 dataset, and the margin increases as sparsity becomes more extreme. Notably, at 99% sparsity, Elsa remains stable ($55.94/40.10$), whereas Wanda+full fine-tuning degrades ($146.37/71.64$) and Wanda+LoRA collapses ($588.3/247.5$). These results show that Elsa remains stable even in this extreme regime, while simple heuristics fail to preserve performance, highlighting the advantage of Elsa ’s more principled approach.
5.3 Realizing benefits with extreme sparsity
Sparsity is practically meaningful only if it yields real deployment gains beyond reducing parameter count. To quantify this, we benchmark end-to-end text generation on LLaMA 2-7B using MACKO (Macko and Boža, 2025), a recent Sparse-Matrix Vector multiplication (SpMV) kernel combined with memory-efficient MACKO format, supporting an acceleration and memory savings for sparse models. Following MACKO’s standard end-to-end protocol, we convert the sparse models produced by Elsa and evaluate on a single NVIDIA RTX 3090, reporting mean latency of token generation (s), throughput (tokens/s), and memory footprint (MB).
Table 1 shows that sparsity yields clear gains in both speed and memory, with improvements becoming pronounced from 70% sparsity onward. At 90% sparsity, Elsa achieves a $2.50\times$ reduction in end-to-end latency and a $2.56\times$ increase in throughput, together with a $4.60\times$ memory reduction compared to the dense baseline. Pushing further to 95% sparsity provides even stronger efficiency gains (up to $4.00\times$ latency reduction, $3.98\times$ throughput increase, and $7.80\times$ memory reduction). These results validate that the high/extreme-sparsity regime targeted by ELSA is not merely an academic frontier: it enables tangible acceleration and substantial memory savings in memory-constrained settings, particularly benefiting the decoding phase where sparse matrix–vector computation dominates.
@@ 5.4 Scaling to large-r models
Figure 5: Perplexity of Gemma-2-27B at 90% sparsity.
We further validate the scalability of our approach by applying Elsa-$_{\text{-L}}$ to 27B-scale (Gemma-2-27B). Specifically, we additionally employ the low-precision optimizer adam8bit (Dettmers et al., 2022) for $x$-update Equation˜7, with Elsa-$_{\text{-L}}$, where we use ($\texttt{bf16},\texttt{fp8}$) for auxiliary state $(u,z)$, respectively. This design reduces the memory footprint of required states (optimizer, auxiliary variables) by 55% compared to Elsa, enabling pruning at 27B scale under limited resources. Additional implementation details can be found in Section˜B.4.
Figure˜5 demonstrates that Elsa-$_{\text{-L}}$ achieves the lowest perplexity among all compared methods, outperforming the strongest competing method by a factor of $4\times$, supporting our main results at scale.
5.5 Cost analysis
To assess effective efficiency (i.e., quality achieved per pruning compute at fixed sparsity), we measure end-to-end pruning cost on identical NVIDIA A100-80GB GPUs and report GPU-hours together with perplexity for LLaMA-2-7B at 90% sparsity (Table 3). One-shot methods (Wanda, SparseGPT) are the cheapest (0.16/0.25 hours with single GPU) but collapse at this sparsity level, yielding unusable perplexities ($\geq 10^{4}$). The best-performing baseline ALPS improves perplexity but still remains far from usable, while requiring substantial wall-clock time (12.57 hours). Overall, these layer-wise pruners are memory-light by design, yet additional pruning compute does not reliably translate into commensurate quality gains under high sparsity.
In contrast, Elsa achieves substantially lower perplexity (26.97/23.14) with a moderate compute budget (7.12 GPU-hours), yielding a markedly better cost–quality point in this regime. Moreover, even when augmented with matched-budget retraining (Wanda+LoRA/Full), one-shot+retrain baselines do not reach comparable quality, indicating that additional compute is more effectively converted into model quality by Elsa at high sparsity.
Table 3: Compute cost vs. perplexity of different pruners on LLaMA-2-7B at 90% sparsity.
Method Wall-clock (h) #GPUs Wiki C4
Wanda 0.159 1 $2.0\times 10^{4}$ $1.0\times 10^{4}$
SparseGPT 0.251 1 1430 864.5
ALPS 12.57 1 248 180.9
Wanda+LoRA 4.03 1 92.66 65.56
Wanda+Full 1.64 4 42.40 34.87
Elsa 1.78 4 26.97 23.14
@@ 6 Discussion
In this work, we confront the problem of moderate sparsity in LLMs through a critical inspection into the current practice, revealing that the prevailing reliance on the sequential layer-wise reconstruction surrogate may have been constraining the path toward more extreme sparsities. This led us to develop Elsa and Elsa-$_{\text{-L}}$, enabling us to push the sparsity from 50–70% up to 80–90% while maintaining strong language modeling performance, where we also observe tangible deployment benefits such as inference acceleration and memory savings, and further showing that our approach remains effective even in more extreme regimes (e.g., 95–99%). Grounding on optimization principles ensures that our principle effectively solves the true LLM objective as is, while also facilitating the development of advanced techniques that are both theoretically sound and effective for sparsifying LLMs, which we believe were instrumental in attaining strong practical results.
@@
There are many promising directions to pursue for future work: (i) alternative efficiency strategies through advanced memory-efficient and derivative-free optimizers, (ii) system-level advancements in memory offloading, and (iii) extensions to advanced architecture such as Mixture-of-Experts (Mu and Lin, 2025) and multi-modal large language models (Yin et al., 2024b). To conclude, our work validates that the frontier of LLM sparsity can still be expanded by offering a concrete strategy supported by strong empirical evidence. We hope it sets the stage for future breakthroughs and innovations in new directions that have thus far received relatively limited attention.
@@ Acknowledgements
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH),RS-2022-II220959, (part2) Few-Shot learning of Causal Inference in Vision and Language for Decision Making), the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) ( RS-2023-00210466, RS-2025-02264052).
@@
Total # datapoints 32768
@@
Sparsity $\lambda$ sched. OPT-125M OPT-1.3B Gemma-2-2B LLaMA-3.2-3B LLaMA-2-7B LLaMA-2-13B
$\eta$ 50% 1e-4 5e-5 5e-5
60% 2e-4 1e-4 1e-4 1e-4
70% 1e-4 5e-5 5e-5
@@
90% 2e-4 1e-4 5e-5 1e-4 1e-4 5e-5
$\lambda$ 50% constant 1e-4 2e-1 1e-3 1e-3 2e-3
60% constant 1e-3 1e-3 2e-3 5e-3 1e-3
70% cosine 2e-3 5e-3 1e-2 5e-3 2e-2
80% cosine 5e-3 2e-2 5e-2
90% cosine 1e-3 1e-3 5e-4 1e-3 1e-3 2e-3
Table 6: Batch size (BS) and number of steps used at different sparsity levels.
Model(s) 50–60% 70–90%
BS Steps BS Steps
OPT-125M 16 2048 8   4096
OPT-1.3B, Gemma-2-2B, LLaMA-3.2-3B, LLaMA-2-7B, LLaMA-2-13B 32 1024